How do you solve the equation, 5/7x+1/3=2/5-2/7x+2/5
Answer by unlockmath(1688)

Hello,
(Note: I'm assuming the equation is like this 5/7(x)+1/3=2/5-2/7(x)+2/5))
5/7(x)+1/3=2/5-2/7(x)+2/5 First add 2/7(x) to both sides and combine 2/5 and 2/5 to get:
7/7(x)+1/3=4/5 Now subtract 1/3 from both sides to get:
x=7/15
There are a couple ways to tackle this problem but this is the quickest.
RJ
Hey, check out a new book I wrote at:
www.math-unlock.com